# The Codecfake Dataset and Countermeasures for the Universally Detection of Deepfake Audio

### v1

@@ Abstract
With the proliferation of Audio Language Model (ALM) based deepfake audio, there is an urgent need for effective detection methods. Unlike traditional deepfake audio generation, which often involves multi-step processes culminating in vocoder usage, ALM directly utilizes neural codec methods to decode discrete codes into audio. Moreover, driven by large-scale data, ALMs exhibit remarkable robustness and versatility, posing a significant challenge to current audio deepfake detection (ADD) models. To effectively detect ALM-based deepfake audio, we focus on the mechanism of the ALM-based audio generation method, the conversion from neural codec to waveform. We initially construct the Codecfake dataset, an open-source large-scale dataset, including two languages, millions of audio samples, and various test conditions, tailored for ALM-based audio detection. Additionally, to achieve universal detection of deepfake audio and tackle domain ascent bias issue of original SAM, we propose the CSAM strategy to learn a domain balanced and generalized minima. Experiment results demonstrate that co-training on Codecfake dataset and vocoded dataset with CSAM strategy yield the lowest average Equal Error Rate (EER) of 0.616% across all test conditions compared to baseline models.
@@ I Introduction
With the development of Audio Language Model (ALM) based audio generation methods, an increasing number of deepfake audios are being produced. Previously, deepfake audio was synthesized through acoustic models and back-end vocoder, categorizing it as vocoder-based audio. In contrast, ALM-based audio generation models adopt only an end-to-end generation approach. They employ the neural codec method to discretize latent representations, transforming the generation task from an autoregressive one into a discrete representation classification task. Unlike vocoders, which commonly require specific datasets for training, ALMs leverage a multitude of diverse audio datasets, facilitating the generated audio with high levels of robustness and versatility. As a result, ALMs not only demonstrate proficiency in synthesizing speech from various speakers but also exhibit their ability to generate audio beyond the realm of speech, encompassing sound effects, singing voices, and music[1]. Currently, the growing prevalence of ALM-based deepfake audio poses significant challenges to current audio deepfake detection (ADD) models.
@@
Figure 2: Mel-spectrogram of the original audio alongside seven codec-based audio samples generated from the original. The top row is generated from VCTK, and the bottom row is generated from AISHELL3.
@@
In this paper, we propose Codecfake, an initial large-scale dataset for detecting ALM-based deepfake audio, encompassing millions of audio samples across two languages. Specifically, we select seven representative open-source neural codec models to construct fake audio. These codec types encompass the current mainstream ALM-based audio generation models. To effectively assess the capability of ADD models trained with Codecfake, we set up diverse test conditions, encompassing unseen codec-based audio, codec-based audio synthesized with varying parameters, and in-the-wild ALM-based audio. For countermeasure, we propose a generalized training strategy called co-training sharpnes aware minimization (CSAM) to effective detect both the vocoder-based audio and codec-based audio. Experiments demonstrate that vocoder-trained ADD models are insufficient for effectively detecting codec-based audio. Leveraging the proposed Codecfake dataset and CSAM strategy for training, we achieve the lowest 0.616% average EER on diverse test condition.
@@ II-A Vocoder-based deepfake audio
In the field of Voice Conversion (VC), the overall architecture of mainstream VC models today is essentially similar to TTS, comprising both an acoustic model and a vocoder. However, the key difference lies in the input to the acoustic model, which is no longer text features but rather specific content features extracted from the audio to be converted. Additionally, speaker characteristics from the target audio are incorporated into the acoustic model. Early content features primarily used Phonetic PosteriorGrams (PPGs) [29], whereas contemporary approaches mainly rely on audio pre-training models such as HuBERT [30] for feature extraction. Speaker characteristics are typically represented by d-vectors [31] or intermediate layer features from speaker verification models.
@@ III-A Overview
Difference. Traditional ADD datasets are primarily constructed using vocoders, while ALM primarily utilizes codec-based methods to decode discrete codes generated by language models into audio, rather than using vocoders for decoding. Therefore, models trained on datasets constructed using vocoders have limited generalization capabilities for detecting ALM-generated fake audio. The main difference between this dataset and previous ones lies in the selection of various neural audio codec models to construct the ADD dataset. Additionally, during training, audio codec models often include data beyond speech, such as general audio, vocals, music, etc., which differs from traditional vocoders that are trained solely on speech. Moreover, audio codec models typically do not use Mel spectrograms as intermediate features during training, which is a notable distinction from traditional vocoders that utilize Mel spectrograms as intermediate features. Training ADD models using a combination of codec-based audio and audio synthesized by traditional vocoders can significantly enhance model generalization. To validate the detection performance of models trained on codec-based audio data for ALM-based audio, we included language-based audio data in the validation set.
@@ III-E Testing Condition
To verify whether codec-trained ADD model can detect audio generated by ALM, we created ALM conditions A1-A3. The generation of these three tasks adopts their open-source default configurations. Among them, A1 and A2 represent text-to-speech (TTS) task, while A3 represents a text-to-audio (TTA) task. A1 is generated by VALL-E [34], A2 by VALL-E X [36], and A3 by AudioGen [33]. For A1 and A2, we selct the text sourced from the VCTK evaluation set corresponding to 4451 text items. However, the speaker p315 does not have corresponding text, so the generated audio for A1 and A2 consists of only 4437 audio samples. For the real source domain of A3, we utilize the training subset of the Audiocaps [65], which was employed during the training phase of AudioGen, generating 49,838 fake audios based on its corresponding texts. Given that some of the original Audiocaps audio links are no longer active, we use the currently available 49,275 audio samples as the real source domain for A3.
@@ IV-A Baseline model
For ADD model, we select three baseline models for evaluation as shown in Figure 4. For the front-end features, we chose the commonly used handcrafted feature in the field of audio deepfake generation, the Mel-spectrogram, to facilitate the analysis of different codec-based audios. On the other hand, we utilized the frozen pre-trained hidden representation of wav2vec2-XLS-R [66] as features. This is because existing research has already convincingly demonstrated the superior performance of wav2vec2-xls-r in the ADD task [67]. The front-end W2V2-XLS-R is pre-trained with 436k hours of unannotated genuine speech data in 128 languages using a contrastive objective. As a result of pre-training, the hidden states from the transformer can represent the genuine speech contextualized information. It is worth mention that we utilize the fifth hidden states from the transformer as the final feature referring to previous frozen hidden states research [68].
For the backbone model, we employ two popular backbone networks in the field of ADD, LCNN and AASIST. LCNN is a lightweight convolution-based backbone featuring an MFM layer designed to effectively filter feature channels conducive to authentication. AASIST represents the state-of-the-art backbone in the ADD domain, introducing a novel heterogeneous stacking graph attention layer that models artifacts across heterogeneous temporal and spectral domains using a heterogeneous attention mechanism and stacked nodes.
@@ VI-D Test on ALM-based audio
In practice, ALM-based audio encompasses not only the reconstruction by neural codec method but also manipulation in content. To assess whether ADD models trained on the codecfake dataset can effectively detect ALM-based audio, we conducted experiments as shown in the Table LABEL:tab:ALM. At first, we observed that employing solely vocoder-trained ADD models yielded satisfactory outcomes for ALM-based audio, particularly evident in evaluations on A1 and A2, where models such as W2V2-AASIST achieved EERs below 1%. However, substantial performance deterioration was evident in evaluations on the condition A3, attributable to OOD issues. The ALM test condition A3 is characterized by sounds with event descriptions that deviate from the distribution of speech. After using Codecfake to train ADD model, performance across A1-A3 are improved, with W2V2-AASIST achieving the lowest average EER of 11.388%, representing a decrease of 4.189% compared to the vocoder-trained ADD models. However, it is noteworthy that while the EER for A3 decreased compared to the vocoder-trained ADD models, it still remains relatively high.
@@ VII Limitation and Future direction
More diverse audio type: Codecfake dataset now incorporates mainstream codec methods and few of ALM-based method in speech. However, the performance of ADD models significantly declines when dealing with non-speech tasks, such as in the A3 scenario. To further develop a universally applicable ADD model, the dataset should not be limited to speech alone. Current methods like ALM can also synthesize other audio types, such as singing voices and audio events. Therefore, ADD dataset needs to include these different types of audio to achieve universally detection.
More acoustic condition: The current synthesized audio in the Codecfake dataset only includes clean conditions, which may lack robustness for detection in complex real-world noise environments. To address this limitation and training a robust ADD model for real-world scenarios, it is necessary to apply various reverberation and noise enhancements to Codecfake dataset.
Generalized ADD methods: The current ADD methods, such as W2V2-AASIST, exhibit a significant performance decline in OOD scenarios. In this paper, we attempted to achieve a generalized ADD model using co-training and CSAM approaches. However, there remains significant for improvement in OOD scenarios such as IWA and A3. Designing a generalized approach that can detect both vocoder-based audio and ALM-based audio is currently a crucial challenge.
Source tracing methods: It is crucial to develop source tracing methods, as this significantly contributes to safeguarding the copyright of ALM developers. Specifically, we need to design a source tracing method with generalization capabilities. This method should be able to perform classification in the ID ALM and identify novel ALM in OOD situations, which can effectively addressing the continuous emergence of ALM algorithms.
@@ VIII Conclusions
Unlike existing vocoder-based deepfake audio datasets, in this paper, we propose the Codecfake dataset for effectively detecting novel ALM-based audio. Codecfake contains seven representative codec synthesis methods, including two languages, millions of audio samples, and various test conditions. Through experiments, we first verified that vocoder-trained ADD models are insufficient to detect the currently novel codec-based audio. Then, ADD models trained on the Codecfake dataset can correctly detect codec-based audio. To effectively detect both vocoder-based audio and codec-based audio, we propose the CSAM training strategy which can optimize both ERM while learning flat minima. Our work for detecting ALM-based audio is preliminary, and some limitations still exist in Section VII. Future endeavors will aim to address these aforementioned limitations.

### v2

@@ Abstract
With the proliferation of Audio Language Model (ALM) based deepfake audio, there is an urgent need for generalized detection methods. ALM-based deepfake audio currently exhibits widespread, high deception, and type versatility, posing a significant challenge to current audio deepfake detection (ADD) models trained solely on vocoded data. To effectively detect ALM-based deepfake audio, we focus on the mechanism of the ALM-based audio generation method, the conversion from neural codec to waveform. We initially construct the Codecfake dataset, an open-source large-scale dataset, including 2 languages, over 1M audio samples, and various test conditions, focus on ALM-based audio detection. As countermeasure, to achieve universal detection of deepfake audio and tackle domain ascent bias issue of original SAM, we propose the CSAM strategy to learn a domain balanced and generalized minima. In our experiments, we first demonstrate that ADD model training with the Codecfake dataset can effectively detects ALM-based audio. Furthermore, our proposed generalization countermeasure yields the lowest average Equal Error Rate (EER) of 0.616% across all test conditions compared to baseline models. The dataset [^1] https://zenodo.org/records/11171708and associated code [^2] https://github.com/xieyuankun/Codecfakeare available online.
@@ I Introduction
With the development of Audio Language Model (ALM) based audio generation methods, an increasing number of deepfake audios are being produced. These ALM-based deepfake audio poses significant challenges to current audio deepfake detection (ADD) methods:
  • Widespread. Compared to the previous generation paradigm that relied on acoustic models and back-end vocoders, the novel ALMs employs a neural codec-based generative paradigm. By integrating language models, it significantly lowers the bar for audio generation, allowing anyone to easily create synthetic content. As a result, ALM deepfake audio is becoming widespread, posing a challenge for current ADD models trained solely on vocoded data.
  • High deception. ALMs can generate high-quality deepfake audio enhancing its deception. They demonstrate robustness by synthesizing high-quality speech even under low-quality and low-resource conditions, posing challenges for current ADD methods.
  • Type versatility. The diversity of audio categories poses challenges for current ADD methods trained by speech. Unlike vocoders, which commonly require specific datasets for training, ALMs leverage a multitude of diverse audio datasets, facilitating the generated audio with versatility type. As a result, ALMs exhibit ability to generate audio beyond the realm of speech, encompassing sound effects, singing voices, and music[1].
@@
Figure 2: Mel-spectrogram of the original audio alongside 7 codec-based audio samples generated from the original. The top row is generated from VCTK, and the bottom row is generated from AISHELL3.
@@
In this paper, we propose Codecfake, an initial large-scale dataset for detecting ALM-based deepfake audio, encompassing over 1M of audio samples across two languages. Specifically, we select 7 representative open-source neural codec models to construct fake audio. These codec types encompass the current mainstream ALM-based audio generation models. To effectively assess the capability of ADD models trained with Codecfake, we set up diverse test conditions, encompassing unseen codec-based audio, codec-based audio synthesized with varying parameters, and in-the-wild ALM-based audio. For countermeasure, we propose a generalized training strategy called co-training sharpnes aware minimization (CSAM) to effective detect both the vocoder-based audio and codec-based audio. Experiments demonstrate that vocoder-trained ADD models are insufficient for effectively detecting codec-based audio. Leveraging the proposed Codecfake dataset and CSAM strategy for training, we achieve the lowest 0.616% average EER on diverse test condition.
@@ II-A Vocoder-based deepfake audio
In the field of Voice Conversion (VC), the overall architecture of mainstream VC models today is essentially similar to TTS, comprising both an acoustic model and a vocoder. However, the key difference lies in the input to the acoustic model, which is no longer text features but rather specific content features extracted from the audio to be converted. Additionally, speaker characteristics from the target audio are incorporated into the acoustic model. Early content features primarily used Phonetic PosteriorGrams (PPGs) [29], whereas contemporary approaches mainly rely on audio pre-training models such as HuBERT [30] for feature extraction. Speaker characteristics are typically represented by d-vectors [31] or intermediate layer features from speaker verification models. In summary, current vocoder-based audio generation methods can produce audio with such high quality and realism that it is often indistinguishable from real audio for people. However, there is still room for improvement in style and diversity.
@@ III-A Overview
Difference. Traditional ADD datasets are primarily constructed using vocoders, while ALM primarily utilizes codec-based methods to decode discrete codes generated by language models into audio, rather than using vocoders. Consequently, models trained on vocoder-based datasets have limited generalization capabilities for detecting ALM-generated fake audio. To address this limitation, this dataset employs various neural audio codec models to construct the ADD dataset. Additionally, during training, audio codec models often include data beyond speech, such as general audio, vocals, music, etc., which differs from traditional vocoders that are trained solely on speech. Moreover, audio codec models typically do not use handcraft acoustic features as intermediate features during training, which is a notable distinction from traditional vocoders that utilize handcraft acoustic features. Training ADD models using a combination of codec-based audio and audio synthesized by traditional vocoders can significantly enhance model generalization. To validate the detection performance of models trained on codec-based audio data for ALM-based audio, we included language-based audio data in the Codecfake Dataset.
@@ III-E Testing Condition
To verify whether codec-trained ADD model can detect audio generated by ALM, we created ALM conditions A1-A3. The generation of these three tasks adopts their open-source default configurations. Among them, A1 and A2 represent text-to-speech (TTS) task, while A3 represents a text-to-audio (TTA) task. A1 is generated by VALL-E [34], A2 by VALL-E X [36], and A3 by AudioGen [33]. For A1 and A2, we selct the text sourced from the VCTK evaluation set corresponding to 4451 text items. However, the speaker p315 does not have corresponding text, so the generated audio for A1 and A2 consists of only 4437 audio samples. For the real source domain of A3, we utilize the training subset of the Audiocaps [65], which was employed during the training phase of AudioGen, generating 49,838 fake audios based on its corresponding texts. Given that some of the original Audiocaps audio links are no longer active, we use the currently available 49,274 audio samples as the real source domain for A3.
@@ IV-A Baseline model
For ADD model, we select three baseline models for evaluation as shown in Figure 4. For the front-end features, we chose the commonly used handcrafted feature in the field of audio deepfake generation, the Mel-spectrogram, to facilitate the analysis of different codec-based audios. On the other hand, we utilized the frozen pre-trained hidden representation of wav2vec2-XLS-R (W2V2-XLS-R) [66] as features. This is because existing research has already convincingly demonstrated the superior performance of wav2vec2-xls-r in the ADD task [67]. The front-end W2V2-XLS-R is pre-trained with 436k hours of unannotated genuine speech data in 128 languages using a contrastive objective. As a result of pre-training, the hidden states from the transformer can represent the genuine speech contextualized information. It is worth mentioning that we utilize the fifth hidden state from the transformer as the final feature, in reference to previous research on frozen hidden states [68], which demonstrated the best performance.
For the backbone model, we employ two popular backbone networks in the field of ADD, LCNN [69] and AASIST [70]. LCNN is a lightweight convolution-based backbone featuring an MFM layer designed to effectively filter feature channels conducive to authentication. AASIST represents the state-of-the-art backbone in the ADD domain, introducing a novel heterogeneous stacking graph attention layer that models artifacts across heterogeneous temporal and spectral domains using a heterogeneous attention mechanism and stacked nodes. Specifically, we employ the backbone of this version of AASIST [71], which utilizes a fully connected layer to down-sample the hidden states of W2V2 from 1024 dimensions to 128, matching the dimension of the front-end of AASIST.
@@ VI-D Test on ALM-based audio
In practice, ALM-based audio encompasses not only the reconstruction by neural codec method but also manipulation in content. To assess whether ADD models trained on the codecfake dataset can effectively detect ALM-based audio, we conducted experiments as shown in the Table VII. At first, we observed that employing solely vocoder-trained ADD models yielded satisfactory outcomes for ALM-based audio, particularly evident in evaluations on A1 and A2, where models such as W2V2-AASIST achieved EERs below 1%. However, substantial performance deterioration was evident in evaluations on the condition A3, attributable to OOD issues. The ALM test condition A3 is characterized by sounds with event descriptions that deviate from the distribution of speech. After using Codecfake to train ADD model, performance across A1-A3 are improved, with W2V2-AASIST achieving the lowest average EER of 11.388%, representing a decrease of 4.189% compared to the vocoder-trained ADD models. However, it is noteworthy that while the EER for A3 decreased compared to the vocoder-trained ADD models, it still remains relatively high.
@@ VII Limitation and Future direction
More diverse audio type. Codecfake dataset now incorporates mainstream codec methods and few of ALM-based method in speech. However, the performance of ADD models significantly declines when dealing with non-speech tasks, such as in the A3 scenario. To further develop a universally applicable ADD model, the dataset should not be limited to speech alone. Current methods like ALM can also synthesize other audio types, such as singing voices and audio events. Therefore, ADD dataset needs to include these different types of audio to achieve universally detection.
More acoustic condition. The current synthesized audio in the Codecfake dataset only includes clean conditions, which may lack robustness for detection in complex real-world noise environments. To address this limitation and training a robust ADD model for real-world scenarios, it is necessary to apply various reverberation and noise enhancements to Codecfake dataset.
Generalized ADD methods. The current ADD methods, such as W2V2-AASIST, exhibit a significant performance decline in OOD scenarios. In this paper, we attempted to achieve a generalized ADD model using co-training and CSAM approaches. However, there remains significant for improvement in OOD scenarios such as IWA and A3. Designing a generalized approach that can detect both vocoder-based audio and ALM-based audio is currently a crucial challenge.
Source tracing methods. It is crucial to develop source tracing methods, as this significantly contributes to safeguarding the copyright of ALM developers. Specifically, we need to design a source tracing method with generalization capabilities. This method should be able to perform classification in the ID ALM and identify novel ALM in OOD situations, which can effectively addressing the continuous emergence of ALM algorithms.
@@ VIII Conclusions
Unlike existing vocoder-based deepfake audio datasets, in this paper, we propose the Codecfake dataset for effectively detecting novel ALM-based audio. Codecfake contains seven representative codec synthesis methods, including two languages, millions of audio samples, and various test conditions. Through experiments, we first verified that vocoder-trained ADD models are insufficient to detect the currently novel codec-based audio. Then, ADD models trained on the Codecfake dataset can correctly detect codec-based audio. To effectively detect both vocoder-based audio and codec-based audio, we propose the CSAM training strategy which can optimize both ERM while learning flat minima. Our work for detecting ALM-based audio is preliminary, and some limitations still exist. Future endeavors will aim to address these aforementioned limitations.